Length of a string tied to two rigid supports is 40 cm.Maximum length (wave length in cm) of a
stationary wave produced on it is
(a) 20 (b) 80 (c) 40 (d)120

Grade:9

1 Answers

Rituraj Tiwari
3 years ago
When the string is plucked in the middle, it vibrates in one loop with nodes at fixed ends and an antinode in the middle.

Length of a string tied to two rigid supports is 40 cm

λ/2 = length

λ = 2 × length = 2 × 40 = 80 cm

Therefore,

Maximum length = 80 cm.